# LLMScan: Causal Scan for LLM Misbehavior Detection

### v2

@@ Abstract
Despite the success of Large Language Models (LLMs) across various fields, their potential to generate untruthful, biased and harmful responses poses significant risks, particularly in critical applications. This highlights the urgent need for systematic methods to detect and prevent such misbehavior. While existing approaches target specific issues such as harmful responses, this work introduces LLMScan, an innovative LLM monitoring technique based on causality analysis, offering a comprehensive solution. LLMScan systematically monitors the inner workings of an LLM through the lens of causal inference, operating on the premise that the LLM’s ‘brain’ behaves differently when misbehaving. By analyzing the causal contributions of the LLM’s input tokens and transformer layers, LLMScan effectively detects misbehavior. Extensive experiments across various tasks and models reveal clear distinctions in the causal distributions between normal behavior and misbehavior, enabling the development of accurate, lightweight detectors for a variety of misbehavior detection tasks.
@@ 1 Introduction
Large language models (LLMs) demonstrate advanced capabilities in mimicking human language and styles for diverse applications (OpenAI, 2023), from literary creation (Yuan et al., 2022) to code generation (Li et al., 2023; Wang et al., 2023b). At the same time, they have shown the potential to misbehave in various ways, raising serious concerns about their use in critical real-world applications. First, LLMs can inadvertently produce untruthful responses, fabricating information that may be plausible but entirely fictitious, thus misleading users or misrepresenting facts (Rawte et al., 2023). Second, LLMs can be exploited for malicious purposes, such as through jailbreak attacks (Liu et al., 2024; Zou et al., 2023b; Zeng et al., 2024), where the model’s safety mechanisms are bypassed to produce harmful outputs. Third, the generation of toxic responses such as insulting or offensive content remains a significant concern (Wang & Chang, 2022). Lastly, biased responses, which can appear as discriminatory or prejudiced remarks, are especially troubling as they have the potential to reinforce stereotypes and undermine societal efforts toward equality and inclusivity (Stanovsky et al., 2019; Zhao et al., 2018).
Numerous attempts have been made to detect LLM misbehavior (Pacchiardi et al., 2023; Robey et al., 2024; Sap et al., 2020; Caselli et al., 2021). However, existing approaches often face two significant limitations. First, they tend to focus on a single type of misbehavior, which reduces the overall effectiveness of each method and requires the integration of multiple systems to comprehensively address the diverse forms of misbehavior. Second, many methods rely on analyzing the model’s responses, which can be inefficient or even ineffective, particularly for longer outputs. Additionally, they are often vulnerable to adaptive adversarial attacks (Sato et al., 2018; Hartvigsen et al., 2022). As a result, there is an urgent need for more general and robust misbehavior detection methods capable of identifying (and mitigating) the full range of LLM misbehavior.
@@
To evaluate the effectiveness of LLMScan, we conduct experiments using four popular LLMs across 13 diverse datasets. The results demonstrate that LLMScan accurately identifies four types of misbehavior, with particularly strong performance in detecting untruthful, toxic responses, and harmful outputs resulting from jailbreak attacks, achieving average AUCs above 0.95. Additionally, we perform ablation studies to evaluate the individual contributions of the causality distribution from prompt tokens and neural layers. The findings reveal that these components are complementary, enhancing the overall effectiveness of LLMScan.
Our contributions are summarized as follows. First, we develop a novel method for “brain-scanning” LLMs through the lens of causality analysis. Secondly, we propose a unified method for detecting various types of misbehavior (including generating untruthful, biased, harmful and toxic responses) based on the brain-scan results. Lastly, we conduct experiments across 52 benchmarks, demonstrating that LLMScan effectively identifies four types of misbehaviors in LLMs. Our code is publicly available at https://github.com/anonymousaa1/LLM_Scan.git.
@@ 2 Background and Related Works
Existing methods for detecting LLM misbehavior have predominantly focused on specific and narrowly defined scenarios (Pacchiardi et al., 2023). While these existing approaches are effective within their specific domains, they often struggle to generalize across different types of misbehavior. In the following, we discuss the need for four examples of LLM misbehavior detection and highlight the necessity for more comprehensive and adaptable detection mechanisms that can be applied across a wide range of misbehavior cases.
Lie Detection: LLMs can “lie”, i.e., producing untruthful statements despite demonstrably “knowing” the truth (Pacchiardi et al., 2023). That is, a response from an LLM is considered a lie if and only if a) the response is untruthful and b) the model “knows” the truthful answer (i.e., the model provides the truthful response under question-answering scrutiny) (Pacchiardi et al., 2023). For example, LLMs might “lie” when instructed to disseminate misinformation.
@@
Jailbreak Detection: Aligned LLMs are designed to adhere to built-in restrictions and ethical safeguards aimed at preventing the generation of harmful or inappropriate content. However, these models can be compromised through the use of cleverly crafted prompts, a technique known as “jailbreaking” (Wei et al., 2024). It has been shown that jailbreaking consistently circumvents the safety alignment of both open-source and black-box models such as GPT-4.
@@
Numerous studies have been conducted to defend against jailbreaking attacks (Robey et al., 2024; Alon & Kamfonas, 2023; Zheng et al., 2024; Li et al., 2024; Hu et al., 2024), which can be broadly categorized into three strategies. The first strategy, prompt detection, seeks to identify and filter out malicious inputs by analyzing their perplexity or matching them against known patterns of jailbreak attempts (Alon & Kamfonas, 2023; Jain et al., 2023). While effective in some cases, these methods are limited by the diversity and variability of jailbreak prompts. The second strategy, input transformation, offers an additional layer of defense by introducing controlled perturbations or alternations to the inputs, such as reordering words or applying noise (Robey et al., 2023; Xie et al., 2023; Zhang et al., 2024; Wei et al., 2023). However, even these methods can be evaded by more sophisticated jailbreak attempts, which has led to the use of behavioral analysis. Lastly, some approaches focus on monitoring the model’s responses for signs of insecure or unexpected behavior, or analyzing internal model states to detect anomalies during inference (Li et al., 2024; Hu et al., 2024).
@@
Toxicity Detection: LLMs may generate toxic content, i.e., responses that contain abusive, aggressive, or other forms of inappropriate content. It is likely because their training data include toxic or inappropriate material, and the models lack the ability to make real-world moral judgments (Ousidhoum et al., 2021). As a result, they struggle to discern appropriate responses without proper contextual guidance. Toxic outputs not only degrade the user experience and increase psychological stress, but they also contribute to broader negative social outcomes, such as the spread of hate speech and the deepening of social divisions.
@@
Existing research on toxic content detection generally follows two main directions. The first focuses on creating benchmark datasets for toxic detection (Vidgen et al., 2021; Hartvigsen et al., 2022). The second involves supervised learning, where models are trained on labeled datasets to identify toxic content. This includes fine-tuning LLMs on toxicity datasets to improve their ability to detect toxic language (Caselli et al., 2021; Kim et al., 2022; Wang & Chang, 2022). However, these supervised learning approaches face several challenges. Firstly, they rely on labeled training data, which is difficult to obtain due to the lack of standardized ground truth. Second, deploying LLMs for toxic content detection in production environments can be computationally expensive, particularly when processing prompts with a large number of tokens.
Bias Detection: Typically trained on vast amounts of internet-based data, LLMs inevitably inherit stereotypes, misrepresentations and derogatory language, and other denigrating behaviors that disproportionately affect already vulnerable and marginalized communities (Touvron et al., 2023b; Biderman et al., 2023). These harms manifest as “social bias”. Extensive efforts and benchmarks have been developed to investigate and mitigate bias and stereotype issues (Stanovsky et al., 2019; Zhao et al., 2018). Those social biases can be categorized into three primary dimensions: 1) generative stereotype bias, which involves using prompts to condition text generation (Dhamala et al., 2021; Parrish et al., 2022); 2) counterfactual bias, where bias is assessed using pairs of sentences that substitute social groups (Nangia et al., 2020; Barikeri et al., 2021; Huang et al., 2020); and 3) representation bias, focusing on the embedding layer (May et al., 2019; Guo & Caliskan, 2021). Each category employs different methods and approaches for detecting biases within text and model representations.
@@ 3.1 Causality Analysis
Causality analysis aims to identify and quantify the presence of causal relationships between events. To conduct causality analysis on machine learning models such as LLMs, we adopt the approach described in (Chattopadhyay et al., 2019; Sun et al., 2022), and treat LLMs as Structured Causal Models (SCM). In this context, the causal effect of a variable, such as an input token or transformer layer within the LLM, is calculated by measuring the difference in outputs under different interventions Rubin (1974). Formally, the causal effect of a given endogenous variable $x$ on output $y$ is measured as follows (Chattopadhyay et al., 2019; Sun et al., 2022):
@@
Note that there are two key reasons for using attention scores for causal effect measurement. First, at the token level, attention scores more effectively capture inter-token relationships, which may be obscured in later results, such as logits. Attention distances provide a clearer reflection of how the model’s understanding shifts with token interventions, offering a more precise measure of each token’s causal impact. Our empirical results reinforce this point, showing that attention head distances, typically ranging between 2 and 3, are more sensitive to logits differences, which generally fell below 0.2. Additionally, previous research has demonstrated that attention layers may encapsulate the model’s knowledge, including harmful or inappropriate information that may require censorship (Meng et al., 2022).
@@
Let $\ell$ be a layer of the LLM; and $x$ be the prompt. The ACE of $\ell$ for prompt $x$ is
@@
Note that here we use the logit as the measure of the causal effect, as the attention is no longer available once a layer is skipped.
The Causal Map. Given a prompt $x$, we systematically calculate the causal effect of each token and each layer to form a causal map, i.e., a heatmap through the lens of causality. Our hypothesis is that such maps would allow us to differentiate misbehavior from normal behaviors. For example, Figure LABEL:fig:heatmap_CE shows two causal maps, where the left one corresponds to a truthful response generated by the LLaMA-2-7B model (with 32 layers), and the right one corresponds to an untruthful response. The prompt used is “Who developed Windows95?”. The truthful response is “Microsoft Corporation”. The untruthful response is “Bill Gates” when the model is induced to lie with the instruction “Answer the following question with a lie”. Note that each causal map consists of two parts: the input token causal effects and the layer causal effects. It can be observed that there are obvious differences between the causal maps. When generating the truthful response, a few specific layers stand out with significantly high causal effects, indicating that these layers play a dominant role in producing the truthful output. However, when generating the untruthful response, a greater number of layers contribute, as demonstrated by relatively uniform high causal effects, potentially weaving together contextual details that enhance the credibility of the lie.
@@ 3.2 Misbehavior Detection
The misbehavior detector is a classifier that takes a causal map as input and predicts whether it indicates certainly misbehavior. For each type of misbehavior, we train the detector using two contrasting sets of causal maps: one representing normal behavior (e.g., those producing truthful responses) and the other containing causal maps of misbehavior (e.g., those producing untruthful responses). In our implementation, we adopt simple yet effective Multi-Layer Perceptron (MLP) trained with the Adam optimizer.
@@
Table 1: Experimental benchmarks.
Task Dataset Model
Lie Detection Questions1000 (Meng et al., 2022)
WikiData (Vrandečić & Krötzsch, 2014)
SciQ (Welbl et al., 2017) Llama-2-(7b/13b) (Touvron et al., 2023a)
CommonSesnseQA (Talmor et al., 2022)
MathQA (Patel et al., 2021)
Jailbreak Detection AutoDAN (Liu et al., 2024)
GCG (Zou et al., 2023b) Llama-3.1 (Dubey et al., 2024)
PAP (Zeng et al., 2024)
Toxicity Detection SocialChem (Forbes et al., 2020)
Bias Detection BBQ-gender (Parrish et al., 2022)
BBQ-religion (Parrish et al., 2022) Mistral (Jiang et al., 2023)
BBQ-race (Parrish et al., 2022)
BBQ-sexualOr (Parrish et al., 2022)
@@ 4 Experimental Evaluation
In this section, we evaluate the effectiveness of LLMScan through multiple experiments. We apply LLMScan to detect the four types of misbehavior discussed in Section 2. It should be clear that LLMScan can be easily extended to other kinds of misbehavior. Table 1 summarizes the 4 tasks, 13 public datasets, and 3 well-known open-source LLMs adopted in our experiments. More details regarding the datasets, their processing, and the models are provided in the supplementary material.
For baseline comparison, we focus on existing misbehavior detectors that rely on analyzing the internal details of the model (rather than the final response). First, for lie detection, we take the approach reported in (Pacchiardi et al., 2023), which is based on the hypothesis that LLM’s behavior would become abnormal after lying. It works by asking a predefined set of unrelated follow-up questions after a suspected lie and feeding the log probabilities of the LLM’s yes/no answers into a logistic regression classifier for lie detection. We selected this approach as the baseline due to its superior detection performance across different LLM architectures and configurations. For the remaining tasks (i.e., jailbreak detection, toxicity detection, and bias detection), most existing approaches rely on analyzing the content of model responses. To ensure a fair comparison, we adopt a recently proposed approach called RepE (Zou et al., 2023a) as a baseline, as it also focuses on analyzing internal model behaviors to monitor LLMs. Specifically, RepE aims to learn a representation of high-level concepts (such as harmlessness) based on internal neural activities. It supports a wide range of detection tasks, including honesty (Lie), emotion, harmlessness (jailbreak), and fairness (bias). Note that we did not use RepE as a baseline for lie detection since it relies on generated false and true statements as input to detect dishonest behaviors. In contrast, our work and the work reported in Pacchiardi et al. (2023) focus on lie-instructed QA scenarios instead of predefined statements, for which the detection task differs fundamentally.
@@
Table 2: Performance (Accuracy and AUC Score) of LLMScan on four misbehavior detection tasks, compared to the baseline (Accuracy). The better results are highlighted in bold.
@@
Task Dataset Llama-2-7b Llama-2-13b Llama-3.1 Mistral
AUC Acc B.Acc AUC Acc B.Acc AUC Acc B.Acc AUC Acc B.Acc
Lie Detection Questions1000 1.00 0.97 0.64 1.00 0.98 0.72 0.97 0.92 0.77 0.99 0.94 0.97
WikiData 1.00 0.99 0.60 1.00 0.99 0.74 1.00 0.98 0.77 1.00 0.97 0.88
SciQ 0.98 0.92 0.69 1.00 0.98 0.74 0.98 0.94 0.82 0.98 0.93 0.96
CommonSenseQA 1.00 0.98 0.70 1.00 1.00 0.63 0.99 0.94 0.68 0.99 0.93 0.88
MathQA 0.93 0.83 0.59 1.00 0.94 0.63 0.92 0.88 0.67 0.97 0.87 0.92
@@
Bias Detection BBQ-gender 0.77 0.72 0.57 0.74 0.69 0.58 0.72 0.74 0.60 0.72 0.77 0.66
BBQ-religion 0.78 0.73 0.58 0.76 0.70 0.57 0.72 0.75 0.60 0.74 0.77 0.67
BBQ-race 0.77 0.70 0.58 0.77 0.71 0.59 0.72 0.72 0.58 0.75 0.79 0.62
BBQ-sexualOr 0.78 0.74 0.56 0.74 0.66 0.55 0.71 0.71 0.60 0.74 0.78 0.66

### v3

@@ Abstract
Despite the success of Large Language Models (LLMs) across various fields, their potential to generate untruthful and harmful responses poses significant risks, particularly in critical applications. This highlights the urgent need for systematic methods to detect and prevent such misbehavior. While existing approaches target specific issues such as harmful responses, this work introduces LLMScan, an innovative LLM monitoring technique based on causality analysis, offering a comprehensive solution. LLMScan systematically monitors the inner workings of an LLM through the lens of causal inference, operating on the premise that the LLM’s ‘brain’ behaves differently when generating harmful or untruthful responses. By analyzing the causal contributions of the LLM’s input tokens and transformer layers, LLMScan effectively detects misbehavior. Extensive experiments across various tasks and models reveal clear distinctions in the causal distributions between normal behavior and misbehavior, enabling the development of accurate, lightweight detectors for a variety of misbehavior detection tasks.
Machine Learning, ICML
[…]
tcb@breakable
@@ 1 Introduction
Large language models (LLMs) demonstrate advanced capabilities in mimicking human language and styles for diverse applications (OpenAI, 2023), from literary creation (Yuan et al., 2022) to code generation (Li et al., 2023; Wang et al., 2023b). At the same time, they have shown the potential to misbehave in various ways, raising serious concerns about their use in critical real-world applications. First, LLMs can inadvertently produce untruthful responses, fabricating information that may be plausible but entirely fictitious, thus misleading users or misrepresenting facts (Rawte et al., 2023). Second, LLMs can be exploited for malicious purposes, such as through jailbreak attacks (Liu et al., 2024; Zou et al., 2023b; Zeng et al., 2024), where the model’s safety mechanisms are bypassed to produce harmful outputs. Third, the generation of toxic responses such as insulting or offensive content remains a significant concern (Wang & Chang, 2022).
Lastly, LLMs are vulnerable to backdoor attacks, where specific triggers in the input prompt cause the model to generate adversarial outputs aligned with an attacker’s goals (Xu et al., 2024; Yan et al., 2024). These attacks can be subtle and hard to detect, as the model’s behavior changes only under certain conditions, making it difficult to distinguish between normal and malicious responses.
Numerous attempts have been made to detect LLM misbehavior (Pacchiardi et al., 2023; Robey et al., 2024; Sap et al., 2020; Caselli et al., 2021), but existing approaches face two major limitations. First, they typically focus on a single type of misbehavior, limiting their effectiveness and requiring multiple systems to address all forms of misbehavior. Second, many methods rely on analyzing model responses, which can be inefficient, especially for longer outputs, and are vulnerable to adaptive adversarial attacks (Sato et al., 2018; Hartvigsen et al., 2022). Therefore, there is a need for more general and robust detection methods that can identify and mitigate a broader range of LLM misbehaviors.
@@
To evaluate the effectiveness of LLMScan, we conduct experiments using four popular LLMs across 13 diverse datasets. The results demonstrate that LLMScan accurately identifies four types of misbehavior, i.e., untruthful, toxic, harmful outputs from jailbreak attacks, as well as harmful responses from backdoor attacks, achieving average AUCs above 0.98. Additionally, we perform ablation studies to evaluate the individual contributions of the causality distribution from prompt tokens and neural layers. The findings reveal that these components are complementary, enhancing the overall effectiveness of LLMScan.
Our contributions are as follows: First, we introduce a novel method for ”brain-scanning” LLMs using causality analysis. Second, we propose a unified approach for detecting various misbehaviors, including untruthful, harmful, and toxic responses, based on these brain-scan results. Lastly, we demonstrate through experiments on 56 benchmarks that LLMScan effectively identifies four types of LLM misbehaviors. Our code is available at https://github.com/zhangmengling/LLMScan.
@@ 2 Background and Related Works
Existing LLM misbehavior detection methods mainly focus on specific scenarios (Pacchiardi et al., 2023). While effective in their domains, they often fail to generalize across different misbehaviors. We highlight four examples of misbehavior detection and the need for more adaptable and comprehensive detection mechanisms in the following.
Lie Detection. LLMs can “lie”, i.e., producing untruthful statements despite demonstrably “knowing” the truth (Pacchiardi et al., 2023). That is, a response from an LLM is considered a lie if and only if a) the response is untruthful and b) the model “knows” the truthful answer (i.e., the model provides the truthful response under question-answering scrutiny) (Pacchiardi et al., 2023). For example, LLMs might “lie” when instructed to disseminate misinformation.
@@
Jailbreak Detection. Aligned LLMs are designed to follow ethical safeguards and prevent harmful content generation but can be compromised through techniques like ”jailbreaking” (Wei et al., 2024a), which bypasses safety measures in both open-source and black-box models, including GPT-4.
@@
Several defense strategies against jailbreaking have been proposed (Robey et al., 2024; Alon & Kamfonas, 2023; Zheng et al., 2024; Li et al., 2024c; Hu et al., 2024), broadly categorized into three approaches. The first, prompt detection, identifies malicious inputs based on perplexity or known jailbreak patterns (Alon & Kamfonas, 2023; Jain et al., 2023), though it struggles with diverse prompts. The second, input transformation, introduces controlled perturbations (e.g., word reordering or noise) to defend against attacks (Robey et al., 2023; Xie et al., 2023; Zhang et al., 2024; Wei et al., 2023), but can still be bypassed by more sophisticated methods. Finally, behavioral analysis and internal model monitoring detect anomalies in model responses or states during inference (Li et al., 2024c; Hu et al., 2024).
@@
Toxicity Detection. LLMs may generate toxic content, such as abusive or aggressive responses, due to training on data that includes inappropriate material and their inability to make real-world moral judgments (Ousidhoum et al., 2021). This leads to difficulties in discerning appropriate responses without contextual guidance, negatively impacting user experience and contributing to social issues like hate speech and division.
@@
Research on toxic content detection has two main approaches: creating benchmark datasets for toxicity detection (Vidgen et al., 2021; Hartvigsen et al., 2022), and supervised learning, where models are trained on labeled datasets to identify toxic language (Caselli et al., 2021; Kim et al., 2022; Wang & Chang, 2022). However, these methods face challenges, including the need for labeled data, which is difficult to obtain, and the high computational cost of deploying LLMs for toxicity detection in production environments.
Backdoor Detection. Generative LLMs are vulnerable to backdoor attacks, where specific triggers in the prompt lead to adversarial outputs (Xu et al., 2024; Yan et al., 2024; Li et al., 2024a; Gu et al., 2017; Hubinger et al., 2024; Li et al., 2024b). For example, the Badnet attack uses the trigger ”BadMagic” to manipulate the output (Gu et al., 2017), while VPI data poisoning targets topics like negative sentiments toward ”OpenAI” (Yan et al., 2024).
Backdoor detection in NLP models falls into two main approaches. The first detects backdoor triggers in input text (Kurita et al., 2020; Sun, ; Wei et al., 2024b), such as the ONION framework, which identifies outlier words (Qi et al., 2021). The second focuses on detecting backdoors within the model itself, even without access to the input (Zhao et al., 2024; Liu et al., 2023), exemplified by the Neural Attention Distillation framework, which mitigates backdoor effects via knowledge distillation (Li et al., 2021).
@@ 3.1 Causality Analysis
Causality analysis aims to identify and quantify the presence of causal relationships between events. To conduct causality analysis on machine learning models such as LLMs, we adopt the approach described in (Chattopadhyay et al., 2019; Sun et al., 2022), and treat LLMs as Structured Causal Models (SCM). In this context, the causal effect of a variable, such as an input token or transformer layer within the LLM, is calculated by measuring the difference in outputs under different interventions (Rubin, 1974). Formally, the causal effect of a given endogenous variable $x$ on output $y$ is measured as follows (Chattopadhyay et al., 2019; Sun et al., 2022):
@@
There are two key reasons for using attention scores to measure causal effects. First, attention scores at the token level capture inter-token relationships more effectively than later outputs, such as logits, which may obscure these connections. Attention distances provide a clearer reflection of how token interventions affect the model’s understanding, offering a more accurate measure of each token’s causal impact. Our empirical results support this, showing that attention head distances, typically between 2 and 3, are more sensitive to logits differences, which generally fall below 0.2. Additionally, prior research has shown that attention layers may encode the model’s knowledge, including harmful or inappropriate content that may need censorship (Meng et al., 2022).
@@
Let $\ell$ be a layer of the LLM; and $x$ be the prompt. The causal effect of $\ell$ for prompt $x$ is
@@
Note that we only consider the logit at the beginning of the first token. Here we use the logit as the measure of the causal effect, as the attention is no longer available once a layer is skipped.
The Causal Map.
Given a prompt $x$, we systematically calculate the causal effect of each token and each layer to form a causal map, i.e., a heatmap through the lens of causality. Our hypothesis is that such maps would allow us to differentiate misbehavior from normal behaviors. For example, Figure LABEL:fig:heatmap_CE shows two causal maps, where the left one corresponds to a truthful response generated by the LLaMA-2-7B model (with 32 layers), and the right one corresponds to an untruthful response. The prompt used is “What is the capital of the Roman Republic?”. The truthful response is “Rome”. The untruthful response is “Paris” when the model is induced to lie with the instruction “Answer the following question with a lie”. Note that each causal map consists of two parts: the input token causal effects and the layer causal effects. It can be observed that there are obvious differences between the causal maps. When generating the truthful response, a few specific layers stand out with significantly high causal effects, indicating that these layers play a dominant role in producing the truthful output. However, when generating the untruthful response, a greater number of layers contribute, as demonstrated by relatively uniform high causal effects, potentially weaving together contextual details that enhance the credibility of the lie. More example causal maps are shown in Appendix D.1.
@@ 3.2 Misbehavior Detection
The misbehavior detector is a classifier that takes a causal map as input and predicts whether it indicates potential misbehavior. For each type of misbehavior, we train the detector using two contrasting sets of causal maps: one representing normal behavior (e.g., those producing truthful responses) and the other containing causal maps of misbehavior (e.g., those producing untruthful responses). In our implementation, we adopt simple yet effective Multi-Layer Perceptron (MLP) trained with the Adam optimizer. More details on the detector settings can be found in Appendix C.2.
@@ 4 Experimental Evaluation
In this section, we evaluate the effectiveness of LLMScan through multiple experiments. We apply LLMScan to detect the four types of misbehavior discussed in Section 2. It should be clear that LLMScan can be easily extended to other kinds of misbehavior. We evaluate LLMScan on 4 tasks: Lie Detection with 5 public datasets (Meng et al., 2022; Vrandečić & Krötzsch, 2014; Welbl et al., 2017; Talmor et al., 2022; Patel et al., 2021), Jailbreak Detection with 3 public datasets (Liu et al., 2024; Zou et al., 2023b; Zeng et al., 2024), Toxicity Detection on dataset SocialChem (Forbes et al., 2020) and Backdoor Detection on datasets generated by 5 different attack methods (Gu et al., 2017; Huang et al., 2024; Li et al., 2024b; Hubinger et al., 2024; Yan et al., 2024). Then, 3 well-known open-source LLMs (Touvron et al., 2023; Dubey et al., 2024; Jiang et al., 2023) are adopted in out experiments. More details regarding the datasets, their processing, and the models are provided in the supplementary material, i.e., Appendix A  B and  C.1.
For baseline comparison, we focus on misbehavior detectors that analyze internal model details rather than final outputs. For lie detection, we consider two baselines. The first, from (Pacchiardi et al., 2023), hypothesizes that LLMs exhibit abnormal behavior after lying, using predefined follow-up questions and the log probabilities of yes/no responses for logistic regression classification. The second, TTPD, detects lies using internal model activations (Bürger et al., 2024). We selected these due to their strong detection performance across various LLMs. For jailbreak and toxicity detection, most methods focus on analyzing model responses. To ensure a fair comparison, we adopt RepE (Zou et al., 2023a) as a baseline, which analyzes internal model behaviors to detect misbehaviors like dishonesty, emotion, and harmlessness. We did not use RepE for lie detection, as it requires both true and false statements, while our work focuses on lie-instructed QA scenarios. For backdoor detection, we use the ONION defense method, which removes potentially outlier tokens from the input (Qi et al., 2021). We define backdoor detection success by the baseline’s ability to mitigate attacks, measured by its detection accuracy.
@@
Table 1: Performance (Accuracy and AUC Score) of LLMScan on four misbehavior detection tasks, compared to accuracy of the baseline (B.Acc). For Lie Detection Task, we consider two baselines, i.e., Baseline 1 is based on Lie Detection (Pacchiardi et al., 2023), and Baseline 2 uses TTPD (Bürger et al., 2024). The results for each baseline are separated by the symbol ‘||||’. The better results are highlighted in bold.
@@
Lie detection Questions1000 1.00 0.97 0.64||||0.97 1.00 0.98 0.72||||0.98 0.97 0.92 0.77||||0.82 0.99 0.94 0.97||||0.92
WikiData 1.00 0.99 0.60||||0.97 1.00 0.99 0.74||||0.99 1.00 0.98 0.77||||0.87 1.00 0.97 0.88||||0.97
SciQ 0.98 0.92 0.69||||0.97 1.00 0.98 0.74||||0.98 0.98 0.94 0.82||||0.84 0.98 0.93 0.96||||0.91
CommonSenseQA 1.00 0.98 0.70||||0.98 1.00 1.00 0.63||||0.99 0.99 0.94 0.68||||0.79 0.99 0.93 0.88||||0.93
MathQA 0.93 0.83 0.59||||0.98 1.00 0.94 0.63||||0.86 0.92 0.88 0.67||||0.77 0.97 0.87 0.92||||0.76
@@
Backdoor Detection Badnet 0.98 0.96 0.48 0.99 0.95 0.71 0.98 0.96 0.44 0.95 0.87 0.54
CTBA 1.00 0.98 0.49 0.99 0.96 0.58 0.99 0.97 0.41 0.99 0.94 0.42
MTBA 0.94 0.89 0.45 0.95 0.89 0.66 0.96 0.91 0.41 0.93 0.88 0.37
Sleeper 0.99 0.98 0.35 0.98 0.96 0.26 0.99 0.97 0.27 0.97 0.91 0.25
VPI 0.97 0.92 0.37 0.99 0.95 0.41 0.98 0.94 0.49 0.95 0.88 0.45